How would I work out the two step equations 13=8-5d?
I added 8 to both sides and then divided both sides by 5 but, that is wrong.
It's what my teacher told me to do.
Do I have to act like 5d is -5d?
Is the equation adding 8 to 5d?

Answer by Taxie(4)

1. Subtract 8 from both sides. You get 13 - 8 = 8 -5d -8. or: 5 = -5d
2. Now divide both sides by - 5. You get. -1 = d
Done.